•   Question 2 of 4  


Question assume ideal and optimal operating conditions for the cycle and the re-heater. Air enters the cycle at 300K and 100kPa and it leaves the combustor at 975K and 1089kPa.

 

Calculate the (a) pressure (kPa) of the air leaving the re-heater



1
Expert's answer
2021-05-24T05:03:02-0400

"\frac{T_b}{T_a}=(\frac{P_b}{P_a})^{\frac{\gamma-1}{\gamma}}"

"\frac{975}{300}=(\frac{1089}{100})^{\frac{\gamma-1}{\gamma}}"

finding gamma

"\frac{975}{300}\cdot \:300=\left(\frac{1089}{100}\right)^{\frac{\gamma-1}{\gamma}}\cdot \:300"

"\gamma=\frac{\ln \left(\frac{1089}{100}\right)}{\ln \left(\frac{1089}{325}\right)}=1.97474"


"\frac{975}{300}=(\frac{P_C}{1089-100})^{\frac{\gamma-1}{\gamma}}"

"\frac{975}{300}=(\frac{P_C}{1089-100})^{\frac{1.97474-1}{1.97474}}"

"\log\frac{975}{300}=(\frac{0.97474}{1.97474})\log\frac{P_c}{989}"

"1.056=\log 
p_c-\log 989"

"\ln log x=\ln 4.0512"

"P_C=11251kP_a"